An arrow is shot upward from a platform of 20ft high with an initial velocity of 184 ft per second. Its height(h)in ft after t seconds is given by: h= 16t squared +184+20. Where will the arrow be 50 ft above the ground?
Found 2 solutions by Alan3354, stanbon:
Answer by Alan3354(69443)

An arrow is shot upward from a platform of 20ft high with an initial velocity of 184 ft per second. Its height(h)in ft after t seconds is given by: h= 16t squared +184+20. Where will the arrow be 50 ft above the ground?
==================
It will be 50 ft above the ground.

Answer by stanbon(75887)

An arrow is shot upward from a platform of 20ft high with an initial velocity of 184 ft per second. Its height(h)in ft after t seconds is given by: h= 16t squared +184+20. When will the arrow be 50 ft above the ground?
-----
h(t) = -16t^2 +184t + 20
------
Solve:: -16t^2 + 184t + 20 = 50
------
-16t^2 + 184t - 30 = 0
------
-8t^2 + 92t - 15 = 0
------
t = 11.3246 seconds
--
Cheers,
Stan H.
=============